### v1

# Could black hole thermodynamics play a role in black hole mergers?

Abstract
Gravitational waves from binary black hole mergers yield values for both the black hole remnant mass $M$ and it’s spin $a$, with the $169$ $a$ values collected so far crowding significantly around their average $\bar{a}=0.6869\pm 0.087$. Could this crowding relate directly to the Davies phase transition point at $a=0.68125$ from black hole thermodynamics? I argue that a necessary challenge for such a connection requires a consistent application of the thermodynamic fluctuation theory that follows from black hole thermodynamics (BHT). Specifically, necessary are a correct choice of fluctuating variables, as well as thermal equilibrium between the event horizon at the Hawking temperature $\sim\mu K$ and the outside universe $\sim 3K$. I show that the former requirement follows in straightforward fashion from the BHT of the Kerr model, while the later requires an accretion disk following the Novikov-Thorne accretion disk model. I construct a thermodynamic fluctuation theory meeting both these requirements. My results open the possibility that black hole mergers are based on some dynamical model (not known to me) with a limiting attractor state at the Davies point.
Keywords: black hole mergers; black hole thermodynamics; thermodynamic fluctuation theory; phase transition; Davies singularity point; Novikov-Thorne accretion disk model;
1 Introduction
The detection of gravitational waves produced by the merger of binary stellar mass black holes (BH) [1, 2, 3, 4] has opened a new era in the study of BH’s. Present methods now allow for the measurement of both the mass $M$ and the spin angular momentum $J$ of the post-merger remnant BH’s. Following the lead of Wu et al. [5], I argue that black hole thermodynamics (BHT) [6, 7] might offer a framework for a striking pattern in the post-merger $(M,J)$ data. Measurements have found that remnant BH’s have dimensionless spin values
$a=\frac{|J|}{M^{2}}\sim 0.7,$ (1)
close to the value of the Davies point (DP) $a=a^{*}=0.68125$ from black hole thermodynamics (BHT) [8]. The DP marks a point where fluctuations in $M$ diverge and thermodynamic stability changes from stable (at $a>a^{*}$) to unstable (at $a<a^{*}$). Is there actually some real causal physical connection between these observed post-merger spin data and the DP? So far as I am aware, this question has not yet been answered. Here I propose to make a good start based on thermodynamic fluctuation theory (TFT).
The DP is naturally grounded in the topic of TFT. To apply this theory here, two major challenges must be overcome. The first challenge is to determine which of the three conserved variables $(M,J,Q)$, with $Q$ the charge, actually fluctuate. A priori, it could be all three, or just one or two, with the other(s) drifting only slowly. A straightforward application of the Kerr BH model [9, 10] resolves this issue. I show that it suffices to consider only $\{M\}$ fluctuations. The second challenge is to establish thermal equilibrium between a stellar mass black hole, with it’s event horizon at the Hawking temperature $(\sim\mu K)$, and the rest of the universe at the cosmic background temperature ($\sim 3$ K). For this, I employ an intermediate co-rotating system to equalize the temperatures.
[…]
Of course, the success that I report here applies only to the observed remnant spin data, because my TFT theory is based on that of just a single Kerr BH. But it is for this observed remnant spin data that we have what appears to be the interesting connection with the DP. Obviously, the full merger process calls for a dynamical theory certainly involving general relativity. But the case put forward here gives a possible context for a full dynamical theory: look for a theory that has the remnant end product of the merger process consisting of the DP as a limiting attractor state. This theory might involve the quantum mechanics that is an essential element of BHT.
2 The observational data
Look first at the observed remnant spin data motivating this paper. These data were taken in four sets of observing runs [1, 2, 3, 4], and are arranged in chronological order in Figure 1, with points colored brown, blue, green, and black, respectively. The observed post-merger spin data $a$ are dimensionless in the system of units given in [8]. On physical grounds, $a$ falls in the range $0\leq a<1$ [8], where $a=0$ corresponds to the non-spinning Schwarzschild black hole, and $a=1$ corresponds to the extremal limit where the black hole spins as fast as is physically possible.
Figure 1: The observed post-merger data: (a) values of the remnant spins $a$ arranged in rough order of increasing time of observation, showing that $a$ tends to crowd around the Davies point at $a=a^{*}=0.68125$, pictured in dashed red, and (b) values of the remnant masses $m$ (in solar units $M_{\odot}$), which show no particular pattern.
Figure 1(a) shows $a$ for the observed remnant spins reported so far. Visually, there is a clear tendency favoring spin values close to the DP, $a=a^{*}=0.68125$. The average observed value of $a$ is $\bar{a}=0.6869\pm 0.087$. There are $98/169$ data points above the DP, in the BHT stable regime, and $71/169$ data points below the DP, in the BHT unstable regime. $151/169$ of the data points ($90\%$) have their error bars encompassing $a^{*}$. There are 10/169 points with error bars completely above the DP, and 8/169 with error bars completely below the DP. These observed remnant spin data make a visual case that the DP might have some relevance to the observed data. For comparison, Figure 1(b) shows the corresponding remnant masses $m$, in solar units. The average value of $m$ is $\bar{m}=57.9\pm 32M_{\odot}$. No particular visual pattern in the distribution of $m$ values is evident.
If BHT requires only that the merger remnants be in the stable thermodynamic regime, then we might expect the data points to skew more above the DP than they actually do. As a measure of this skew, define the skew ratio
[…]
For the observed remnant spin data, $sr=0.58$. If the remnants were always BHT stable, then we might expect future remnant data to tend towards $sr\to 1$. If, on the other hand, the actual clustering of values of $a$ were distributed randomly about the DP then we would expect to find roughly as many values below the DP as above, and $sr$ would tend towards $0.5$, close to where they are now.
Ref. [1] mentions this crowding of the observed $a$ values around the value 0.7, saying “the medians of all final spin distributions are around approximately 0.70.7.” A good source of recent BH discoveries, including spin values around $a=0.7$, is the blog by Carlisle [12].
The Appendix adds some additional statistical observations.
3 The relation to thermodynamic stability
Wu et al. [5] brought in the DP, and presented a statistical analysis of the first three sets of observing runs [1, 2, 3], concluding that, with high probability, “thermodynamic stability is robust for $83$ black hole remnants.” Here, I have added a fourth observing run [4] (the largest) to the analysis, and did not get quite the skew towards $a>a^{*}$ that Wu et al. reported. While the $a$ values in the observed remnant spin data clearly tend to cluster around $a^{*}$, the points show no overwhelming tendency to pick the stable phase over the unstable one. Nevertheless, thermodynamic stability is essential for BHT, and maybe for real BH’s, so I start my discussion of BHT with the theme of stability.
Wu et al. [5] employed the Kerr black hole model [10] in their analysis. Kerr is a general relativistic, spinning, uncharged, and asymptotically flat model parameterized by $(M,J)$. Davies [8] pointed out that while stellar mass BH’s generally have slow spin down rates, the electric discharge rate to $Q=0$ is very fast. Hence, the chargeless Kerr model seems an excellent try at modeling real BH’s.
The Bekenstein-Hawking equation sets the BHT entropy $S=S(M,J)$ to the area $A$ of the event horizon. For Kerr, Davies [8] wrote:
$S(M,J)=\frac{1}{4}\left(M^{2}+\sqrt{M^{4}-J^{2}}\right).$ (3)
I use two systems of units in this paper: geometric units with $M$, $J$, and $S$ expressed in meters, meters squared, and meters squared, respectively, and standard MKS units with lower case mass $m$, angular momentum $j$, and entropy $s$ with [13]
$m=\left(\frac{c^{2}}{G}\right)M,$ (4)
and
$j=\left(\frac{c^{3}}{G}\right)J.$ (5)
Here, $c$ is the speed of light, and $G$ is the gravitational constant. To convert the entropy to real units, we have
$\frac{s}{k_{B}}=S\left(\frac{8\pi}{L_{p}^{2}}\right),$ (6)
where $k_{B}$ is Boltzmann’s constant, and $L_{p}$ is the Planck length
$L_{p}=\sqrt{\frac{\hbar G}{c^{3}}},$ (7)
with $\hbar$ Planck’s constant divided by $2\pi$. The capitalized symbols $G$ and the solar mass $M_{\odot}$ are in MKS units. To be below the extremal value of rotation, $J$ must lie in the range $(-M^{2},M^{2})$. As emphasized by Bekenstein [14], a theory containing both $\hbar$ and $G$ has a quantum gravity flavor.
The BHT temperature $T$ is given by
$\frac{1}{T}=\left(\frac{\partial S}{\partial M}\right)_{J}=\frac{\left(1+\sqrt{1-\frac{J^{2}}{M^{4}}}\right)M}{2\sqrt{1-\frac{J^{2}}{M^{4}}}}.$ (8)
Note that $T\to 0$ as $a\to 1$. This marks the extremal limit, where the BH has its maximum spin rate.
The heat capacity $C_{J}$ at constant $J$ is given by:
$\frac{C_{J}}{M^{2}}=\frac{T}{M^{2}}\left(\frac{\partial S}{\partial T}\right)_{J}=\frac{1-a^{2}+\sqrt{1-a^{2}}}{2+2a^{2}-4\sqrt{1-a^{2}}},$ (9)
where I have used the identity
$\left(\frac{\partial S}{\partial T}\right)_{J}=\left(\frac{\partial S}{\partial M}\right)_{J}/\left(\frac{\partial T}{\partial M}\right)_{J}.$ (10)
With Eq. (9) we may show that $|C_{J}|$ diverges at the DP, is negative at all $a<a^{*}$, and positive at all $a>a^{*}$; see Figure 2. Wu et al. termed these negative and positive sign regimes as thermodynamically unstable and stable, respectively, following textbook terminology [15, 16].
Figure 2: The heat capacity as a function of the spin $a=|J|/M^{2}$. $C_{J}$ diverges and changes sign at the Davies point $a=a^{*}=0.68125$, indicated with the dashed red line.
4 Basic thermodynamic fluctuation theory
Thermodynamic stability is usually presented in statistical mechanics textbooks [15, 16] in the context of thermodynamic fluctuation theory (TFT). TFT requires thermodynamic stability for its application. But, a priori this certainly does not force BH’s to live only in thermodynamically stable states. One can easily imagine that a BH could be held together by its powerful gravity regardless of its BHT stability. In this section, I present the basics of TFT, with new elements specific to this discussion added in subsequent sections.
Consider two thermodynamic systems $A$ and $B$. The composite $A$, $B$ system is thermodynamically stable if a steady state equilibrium persists between the two systems, up to small fluctuations about the state with maximum entropy. Pick $A$ as the black hole and $B$ as its environment, which is here the whole rest of the universe. Mediated by Hawking radiation [17], $A$ and $B$ are expected to exchange certain of their conserved parameters. The fluctuating parameters for $A$ are picked from the set $(M,J,Q)$. This basic scenario is shown in Figure 3(a).
Refer to caption Refer to caption
Figure 3: Sketch of two basic structures framing thermodynamic fluctuation theory for black holes: (a) a spinning black hole surrounded by a non-spinning environment (the rest of the universe), and (b) a spinning black hole connected to the non-spinning environment via an intermediate spinning accretion disc. Fluctuations in the black hole are mediated via Hawking radiation [17] at the event horizon, indicated with a red circle. I start my discussion of the Davies point with scenario (a), and then argue that a correct picture must include the accretion disc in (b) in order to achieve thermal equilibrium at the event horizon.
Consider the possibility that one or two of the variables $(M,J,Q)$ fluctuate much more slowly than the other one or two active fluctuating variables, and be effectively fixed in comparison. This leads to seven possible fluctuation cases: fluctuating $\{M,J,Q\}$, $\{J,Q\}$, $\{M,Q\}$, $\{M,J\}$, $\{M\}$, $\{J\}$, and $\{Q\}$. Which one of these scenarios nature picks at the event horizon depends on the detailed dynamics of Hawking radiation, which I am not sure has been worked out to completion. However, if we require that our thermodynamic fluctuation formulas include a DP, necessary to connect to the observed post-merger spin data, then which variables fluctuate clarifies.
But before we resolve this issue, let me explain the basic structure in TFT. For our two system universe $U=A+B$, the system entropies add to give the entropy of the universe:
$S^{U}(X^{A},X^{B})=S^{A}(X^{A})+S^{B}(X^{B}).$ (11)
Here, $X^{U}$, $X^{A}$, and $X^{B}$ are the lists of values of the conserved fluctuating parameters of the systems in play here, including the whole universe $U$. The entropies of $U$, AA, and $B$ are denoted by $S^{U}(X^{A},X^{B})$, $S^{A}(X^{A})$, and $S^{B}(X^{B})$, respectively. The conservation laws yield $X^{U}=X^{A}+X^{B}$, which leads to
$X^{B}=X^{U}-X^{A},$ (12)
with $X^{U}$ a list of constants. We may then write $S^{U}$ as a function of just $X^{A}$:
$S^{U}(X^{A})=S^{A}(X^{A})+S^{B}(X^{U}-X^{A}).$ (13)
I first take all three parameters $X^{A}=(M,J,Q)$ to be fluctuating, with straightforward generalization to fixing one or two parameters given later. Einstein’s famous thermodynamic fluctuation equation is [15, 16, 18]
$P(X^{A})\propto\exp\left[\frac{S^{U}(X^{A})}{k_{B}}\right]$ (14)
for the probability density $P(X^{A})$ of finding the thermodynamic state of $A$ in some phase space volume element $dMdJdQ$ located at coordinates $X^{A}$. This fluctuation formula was written in the black hole scenario by Åman, Bengtsson, and Pidokrajt [19]. To keep it simple, I use geometric units in this section. In physical applications, we must convert $S^{U}$ in Eq. (14) to real units. But I avoid this here to reduce clutter. In Section $8$, I show an application with real units.
Assuming thermodynamic stability, the equilibrium state has maximum total entropy = $S_{0}$, and coordinate values $X^{A}=X^{A}_{0}$. Expanding $S^{U}(X^{A})$ to second-order about $X^{A}_{0}$ in powers of $\Delta X^{A}=(X^{A}-X^{A}_{0})$, yields
$S^{U}(X^{A})=S_{0}+\frac{\partial S^{A}}{\partial X^{A}_{i}}\Delta X^{A}_{i}-\frac{\partial S^{B}}{\partial X^{B}_{i}}\Delta X^{A}_{i}+\frac{1}{2}\frac{\partial^{2}S^{A}}{\partial X^{A}_{i}\partial X^{A}_{j}}\Delta X^{A}_{i}\Delta X^{A}_{j}+\frac{1}{2}\frac{\partial^{2}S^{B}}{\partial X^{B}_{i}\partial X^{B}_{j}}\Delta X^{A}_{i}\Delta X^{A}_{j}+\cdots,$ (15)
where we sum repeated indices $i$ and $j$ over the parameter list $X^{A}$. The conservation laws yield $\Delta X^{B}=-\Delta X^{A}$, and all the derivatives are evaluated at maximum total entropy.
In this expansion, $S_{0}$ may be absorbed in the normalization constant implicit in Eq. (14). For the maximum $S^{U}(X^{A})$, the first-order terms in the expansion must add to zero, requiring the equality between the corresponding conjugate parameters of $A$ and $B$:
$\frac{\partial S^{A}}{\partial X^{A}_{i}}=\frac{\partial S^{B}}{\partial X^{B}_{i}}.$ (16)
Generally, we have [13]
$\left(\frac{\partial S}{\partial M}\right)_{J,Q}=\frac{1}{T},\,\,\,\left(\frac{\partial S}{\partial J}\right)_{M,Q}=-\frac{\Omega}{T},\,\,\,\mbox{and}\left(\frac{\partial S}{\partial Q}\right)_{M,J}=-\frac{\Phi}{T},$ (17)
where $\Omega$ is the angular velocity, and $\Phi$ is the surface electric potential. At the maximum entropy, the second-order terms must add together to a negative number.
Assume now that the environment $B$ has some ordinary thermodynamics (details unknown and not needed) where both $X^{B}$ and $S^{B}(X^{B})$ scale up linearly with the volume $V^{B}$ of $B$. The Hessian of $S^{B}$ in Eq. (15) now has the scaling factor $V^{B}$ one time in it’s numerator, and twice in it’s denominator. So, as $V_{B}$ increases, this Hessian goes to zero no matter what the function $S^{B}(X^{B})$ is, while the Hessian of $S^{A}$ (with $S^{A}$ given by the Kerr metric) remains fixed. Hence, with $V^{B}\gg V^{A}$, we need only include the known Hessian of $S^{A}$ in Eq. (15), and we have to second order:
$S^{U}(X)=S_{0}+\frac{1}{2}\frac{\partial^{2}S}{\partial X_{i}\partial X^{j}}\Delta X_{i}\Delta X_{j},$ (18)
where we have dropped the superscript $A$ and have the variables $X_{i}$ be the fluctuating BH parameters. If this Hessian matrix is negative definite, then we will have thermodynamic stability.
5 Which conserved parameters actually fluctuate?
Which conserved parameters in the list $(M,J,Q)$ actually fluctuate? A dynamical study of Hawking radiation could be tried here, but there are difficult unknowns. For example, are dark matter particles involved in the fluctuations? But the fluctuating variables are easily sorted out with just thermodynamic fluctuation theory, using no microscopic dynamics at all. Work with the Kerr or the Kerr-Newman model for BH’s, and search the seven possible candidate cases: $\{M,J,Q\}$, $\{J,Q\}$, $\{M,Q\}$, $\{M,J\}$, $\{M\}$, $\{J\}$, and $\{Q\}$ for a DP. The cases with a DP are suitable for making some sense of the observed remnant spin data, and the cases without one are unsuitable.
Ruppeiner [13] checked all the seven fluctuation cases. Because BH’s shed excess charge quickly [8, 20], we need consider only cases with the average $Q$ fluctuations set to zero. Here is a summary of the results: $\{M,J,Q\}$ fluctuations have no regime of stability, and no DP. They are unsuitable. $\{J,Q\}$ fluctuations are stable everywhere, but have no DP. They are unsuitable. $\{M,Q\}$ and $\{M\}$ fluctuations each have DP’s, and are stable for $a>a^{*}$, and unstable for $a<a^{*}$. They are both suitable. $\{M,J\}$ fluctuations have no regime of stability and no DP. They are unsuitable. $\{J\}$ fluctuations are stable everywhere, but have no DP. They are unsuitable. $Q$ fluctuations are stable everywhere, but have no DP. They are unsuitable.
So, all four cases with fluctuating $J$ are unsuitable. Dynamically, this may be because $J$ is generally slow developing [8]. The purely $\{Q\}$ fluctuations are unsuitable. This leaves $\{M\}$ and $\{M,Q\}$ fluctuations as the only suitable cases. But in the $\{M,Q\}$ fluctuations, the fluctuations in $M$ are the same as those in the $\{M\}$ fluctuations, and the fluctuations in $Q$ tend to be tiny, with a maximum standard deviation $\sqrt{\langle(\Delta Q)^{2}\rangle}=3.301\,e$ at $J=0$. Here, $e$ is the electron charge. Hence, we need only work with $\{M\}$ fluctuations in this paper.
I add that Sahay [21] also gave a general discussion of constrained fluctuations of the type above.
6 A co-rotating environment is required for thermal equilibrium
[…]
Approaching this thermal equilibrium problem is relatively simple in its basic structure; we introduce an accretion disk located between the BH and the rest of the outside universe. Let system $A$ be the BH, let $B$ be an intermediate system between $A$ and $C$, and let $C$ be the non-rotating rest of the universe; see Fig. 3(b). Rotational drag has us expect $B$ to acquire some measure of co-rotation with $A$. We imagine that the part of $B$ in direct contact with $A$ could now achieve thermal equilibrium with $A$, causing the first order terms in Eq. (15) to cancel for the $\{M\}$ fluctuation case. $S^{A}$ is likely to be much more massive than $S^{B}$ because $A$ will always contain hugely more matter than $B$. This makes the Hessian of $S^{A}$ in Eq. (15) much larger than that of $S^{B}$, and so we could drop the $S^{B}$ Hessian in Eq. (15), leaving us with the standard thermodynamic fluctuation formula Eq. (18). There seems no need to assume anything about any interaction between $B$ and $C$ as the interface between these two systems has little to do with $A$.
The structure of an accretion disk presents a considerable theoretical challenge. Note that accretion discs generally become intensely hot from in-falling material. In the centers of distant young galaxies, accretion discs around supermassive black holes are the brightest objects (quasars) in the universe, making thermal equilibrium with the cool event horizon seem very unlikely. But a closer look suggests otherwise. A productive accretion disk theory is the Novikov-Thorne disk model (NT) [22]. A theme addressed by this theory is the innermost stable circular particle orbit (ISCO) in the accretion disc. ISCO has radius $r_{isco}$, and lies outside the outer event horizon. $r_{isco}$ depends on $M$, $a$ and the particle orbital direction, prograde or retrograde. Circular orbits with radii close to, but outside $r_{isco}$, have angle $\pi/2$ to the BH rotational axis. The disc of such orbits is very thin and rapidly rotating.
A key finding of NT is that as $r\downarrow r_{isco}$, the internal stresses, the emitted heat, and the temperature are each predicted to go to zero; see Equation (11) of Reynolds [20]. Inside the ISCO, orbits cease to be stable and are predicted to plunge rapidly to the event horizon because the centrifugal force is no longer strong enough to resist gravity. There seems to be no consensus about the temperature between the ISCO and the event horizon, and this region is not physically observable. Nevertheless, there certainly seems a possibility that the temperature there could be low enough to equilibrate with the Hawking radiation. At present, we can only assume this, and we must, since else we can have no thermodynamic equilibrium, and TFT between system $A$ and $B$ fails, leaving us with no connection between the DP and the observational merger data other than coincidence.
7 Have unstable black holes been found?
Do thermodynamically unstable BH’s with $a<a^{*}$ exist, or does BHT rapidly transform unstable BH’s into one or more stable BH’s? Certainly in ordinary thermodynamics, like van der Waals, unstable thermodynamic states do not generally last very long. But the BH scenario could be different. Consider the unstable Schwarzschild BH ($a=0$). Schwarzschild has been a mainstay of BH theory for well over a century. The notion that it actually could not physically exist would probably strike theorists as very unlikely. One would expect its powerful gravity to hold it together. Much more discussed for stellar mass Schwarzschild is the possibility of extremely slow evaporation by Hawking radiation.
However, have unstable BH’s actually been discovered? It is possible that they have. Of the observed post-merger spin data, 8/169 ($\sim 5\%$) have upper and lower error bars both below the DP, with the lowest spin at ${a}=0.14\pm 0.06$. Of course, this unstable group is a minority, and it is conceivably spurious. Certainly, new measurements with smaller error bars might change this picture.
Another arena to search for unstable, slowly spinning BH’s is in pre-merger spin studies [1, 2, 3, 4], which employ for analysis a single spin related parameter $\chi_{\mbox{\tiny{eff}}}$, defined in Equation (4) of [1]. But $\chi_{\mbox{\tiny{eff}}}$ mixes in the spins of both BH’s together with their orbital angular momenta. Separating out the individual spins seems difficult.
Yet another source of BH spin measurements is supermassive BH’s interacting with their accretion discs. This interaction can be probed with X-ray reflection spectroscopy and with thermal continuum fitting. Reynolds [20] reviewed the large literature of measurements of $a$ in these settled supermassive BH’s. Their Table $1$, plotted in their Figure 6, shows $32$ cases with known BH masses ranging from $1.1\times 10^{6}M_{\odot}$ to $4,500\times 10^{6}M_{\odot}$. I replot these results in my Figure 4. The majority of these cases show large spins, many very near the extremal limit $a=1$, and all have their error bars for $a$ either above or encompassing the Davis point.
Figure 4: The spin $a$ as a function of the black hole mass $m$ (in units of $10^{6}$$M_{\odot}$) for supermassive black holes [20]. The blue dots have their upper error bars at $a<1$, and the red dots have their upper errors bars at $a=1$, with their data points placed at their lower error bars.
The supermassive BH data is in sharp contrast to the observed remnant spin data. A count of its points shows that their skew ratio $sr=27/32=0.84$. If we put all the red points in Fig. 4 in the middle of their error bars, rather than at the bottoms, then we get $sr=30/32=0.94$. These $sr$ values support a stability picture with $sr\to 1$, rather than a purely clustering picture with $sr\to 1/2$.
We could be of two minds about this supermassive BH data, shown in Fig. 4. We could argue that this is support for a requirement that all BH’s must be BHT stable. Alternatively, we could argue that supermassive BH’s are naturally spun up very fast by in-falling matter, and that because of their huge masses, they simply spin down very slowly. This would have nothing to do with BHT. Regardless, the supermassive BH data contrasts sharply with the observed post-merger spin data. This is no surprise because the formation pictures of these respective objects probably differ substantially, as do their ages, billions of years on the one hand, and roughly a single second on the other.
Both recently merged and settled black holes reveal that fast spins in the BHT stable regime are the norm. For researchers looking to make a case that BH’s must have their BHT’s in the stable regime, these data would offer support.
8 Are black hole thermodynamic fluctuations physically relevant?
Turn now to BHT mass fluctuations, which diverge at the DP. A stable BH, with $a>a^{*}$, spinning down because of various physical drag processes will inevitably encounter the DP at $a=a^{*}$. Might the BH disintegrate there from the fluctuation instability? If so, could this potentially help to explain the apparent relative sparsity of BH’s in the unstable regime? Below, I write down the BHT mass fluctuation formula, but with no evidence for any related BH disintegrations.
Working out fluctuation formulas in MKS units starting from expressions in geometric units requires care. Here are the steps to follow for $\{M\}$ fluctuations: 1) Start with Eq. (14) in the standard form
$P\propto\exp\left[\frac{\Delta s}{k_{B}}\right],$ (19)
where $\Delta s$ is the deviation of the entropy (in MKS units) from its maximum value at equilibrium. $\Delta s$ is related to the corresponding $\Delta S$ by the conversion formula Eq. (6). Here $\Delta S$ denotes $(S^{U}(X)-S_{0})$ in Eq. (18). $S$ is given by the Kerr equation Eq. (3). 2) Simple algebra yields the Gaussian expression for the probability distribution for $M$:
$P(M)dM=c\exp\left[-\frac{1}{2}g\,\Delta M^{2}\right]dM,$ (20)
where
$g=-\left(\frac{8\pi}{L_{p}^{2}}\right)\frac{d^{2}S}{dM^{2}},$ (21)
and $c$ is a normalization constant. 3) Gaussian integrations yield $c=\sqrt{g/(2\pi)}$ and the second fluctuation moment:
$\langle(\Delta M)^{2}\rangle=\frac{1}{g}.$ (22)
4) Evaluate $g$ using the Kerr formula Eq. (3) for $S$ and make the replacement $J\to aM^{2}$. 5) To switch to MKS units, multiply twice by the conversion factor in Eq. (4). For $a>a^{*}$, this yields the exact expression [13] [^3] There is a small typo in Table II of ref. [13]. In Table II for $\{M\}$ fluctuations with $Q=0$, the term $(1+K)^{2}$ on the right hand side should be just $(1+K)$, as it is in Table I for $\{M,Q\}$ fluctuations with $Q=0$.
$\sqrt{\langle(\Delta m)^{2}\rangle}=\displaystyle 6.1412\times 10^{-9}\,\mbox{kg}\,\sqrt{\frac{\left(1-a^{2}\right)^{3/2}}{a^{2}\left(\sqrt{1-a^{2}}+3\right)-\sqrt{1-a^{2}}-1}}.$ (23)
$\{M,Q\}$ fluctuations with $Q=0$ and $a>a^{*}$ have the same $m$ fluctuations as given in Eq. (23) as well as tiny $q$ fluctuations: $\sqrt{\langle(\Delta q)^{2}\rangle}\leq 3.301e$. These tiny $q$ fluctuations would seem to have little physical effect, so I will just neglect them and work only with the $\{M\}$ fluctuations. Figure 5 shows these $m$ fluctuations, and their divergence as $a\downarrow a^{*}$. A nice asymptotic approximation for Eq. (23) is
$\sqrt{\langle(\Delta m)^{2}\rangle}=\frac{1.6278\times 10^{-9}}{\sqrt{a-a^{*}}}\mbox{kg}.$ (24)
Figure 5: The fluctuations $\sqrt{\langle(\Delta m)^{2}\rangle}$ (in solar mass units $M_{\odot}$) as a function of $(a-a^{*})$. The exact and the asymptotic expressions are nearly identical.
Clearly, the agreement in Fig. 5 between the exact and the asymptotic expression is excellent, and show a power law divergence $\propto(a-a^{*})^{-1/2}$ at the DP. As we can see, for the fluctuations in $m$ to reach an appreciable fraction of a solar mass requires a tiny value of $(a-a^{*})$, so small as to make one wonder if these large mass fluctuations are physically accessible at all.
9 Conclusion and future directions
In conclusion, gravitational wave observations have measured the spins $a$ of newly formed remnant stellar mass black holes (BH) resulting from the mergers of binary black hole systems. These spin values are found to cluster around the Davies singularity point (DP), which has $a=a^{*}=0.68125$. This suggests a significant connection between the results from gravity wave measurements and black hole thermodynamics (BHT).
This paper emphasizes this numerical connection, including the latest (fourth) observed remnant spin data set. I have also contrasted the observed remnant spin data with that of the spin data from much older supermassive black holes. The later tend to spin very fast, with their BHT states overwhelmingly in the stable regime above the DP, and no clustering of spin values about any particular number.
Most significantly, I have argued that the DP has its physical roots in thermodynamic fluctuation theory (TFT). In this context, the challenge is to construct a picture with the correct choice of fluctuating variables, as well as thermal equilibrium at the event horizons. I have met this challenge by arguing that in the Kerr model scenario we need consider only $\{M\}$ fluctuations, and proposing that thermal equilibrium at the event horizon could be achieved with an accretion disk following the Novikov-Thorne disk model (NT). I emphasize that if my approach fails, and is not replaced by a better one, we will be left with no connection between the data clustering in evidence here and BHT. In this case, the correlation shown in Fig. 1(a) will be mere coincidence, a disappointing outcome.
Of course, just establishing the relevance of the DP in this scenario still leaves us short of an actual dynamical mechanism by which the post-merger remnants form. General relativity seems to be the place to start looking for such dynamics. But the case put forward here suggests that it might be productive to try to build some dynamical theory that contains the DP as an attractor state for the final merged remnant black hole. This adds quantum mechanics to the theoretical mix with general relativity as a try to understanding BH mergers.
Finally, I list some future research projects suggested by this paper:
First, to continue to collect BH merger remnant data and see how $sr$ in Eq. (2) develops as the data collection builds.
Second, to find some microscopic justification for the BHT’s findings about the fluctuating variables: the domination of the $M$ fluctuations, the tiny $Q$ fluctuations, and the absence of a DP in the $J$ fluctuations. In these fluctuations, what kind of particles are involved? Does dark matter play a role?
Third, a further consideration of thermal equilibrium at the event horizon, an essential theme in my approach. Can the environment just outside the event horizon really get cold enough to match the Hawking temperature?
Finally, have any astronomical events been seen that might match a BH exploding as it spins down through the DP? Could a BH even get close enough to the DP for fluctuations in $M$ to get large enough to generate sufficient fluctuations? Or are the dominant fluctuating particles too granular to make this likely?
10 Appendix
In this Appendix, I present some additional statistical details that further a case that the observed remnant spin data $a$ cluster tightly around the DP. Let $N_{>}(a)$ be the number of observed remnant spin values above some general $a$. $N_{>}(a)$ is plotted against $a$ in Figure 6. In the data tables that I consulted [1, 2, 3, 4], $a$ is expressed as two digit numbers $0.00$, $0.01$, $0.02$, $\ldots 0.99$, $1.00$, and this gives my figure a somewhat down-stepping look. The red cross near the middle of Figure 6(b) has its horizontal bar set to half the number of particles, and its vertical bar set to $a^{*}$. Notice that the cross center falls close to the data curve. But the uncertainty in $\bar{a}$ is large enough that the cross center could, within error, fall on the data curve itself.
The following possible scenario is then consistent with the present observed remnant spin data: the BH remnants emerge from the mergers with spins very near $a^{*}$, much closer than the observed error bars in $a$. These error bars prevent us from directly seeing such a tight clustering. Regardless the issue of whether the BH’s are in the stable regime or not, we might expect that the observed values of $a$ are roughly equally split between above and below the DP, and with more data points a skew ratio near $sr=1/2$ would become more pronounced.
Refer to caption Refer to caption
Figure 6: $N_{>}$ plotted as a function of $a$, where $N_{>}$ is the number of observed remnant spin data points with spins larger than $a$: a) has all $169$ points included, and b) is focused on the DP. The red cross is centered at $a^{*}=0.68125$ and $N_{>}=169/2$.
Finally, I evaluate $\chi^{2}$, a standard measure of the quality of a fit between theory and data:
$\chi^{2}=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(a_{t}-a_{i})}{\sigma_{i}}\right]^{2},$ (25)
where $n$ is the number of data points, $a_{i}$ is the value of $a$ for the $i$’th data point, $\sigma_{i}$ is the length of the $i$’th error bar, and $a_{t}$ is the theoretical value for $a$ being tested. The values $a_{i}$ generally have different upper and lower error bars. To select between them, I set $\sigma_{i}$ to the upper error bar if $a_{i}$ is below $a_{t}$, and I set $\sigma_{i}$ to the lower error bar if $a_{i}$ is above $a_{t}$. The best fitting value of $a_{t}$ to the data is the one that minimizes $\chi^{2}$.
The full data set in play in Fig. 1(a) is a bit problematic for $\chi^{2}$ analysis. There are a number of outlier points having relatively small error bars, and deviating widely from the pack of points clustered around the DP. But for $a_{t}$ near $a^{*}$ (where $\chi^{2}$ turns out to be smallest) these outlier points contribute unreasonably large amounts to the sum, and run the danger of skewing the analysis unreasonably. Hence, for the $\chi^{2}$ analysis, I omitted the $18$ points whose error bars do not encompass $a^{*}$, leaving $151$ points in the data set. Figure 7 shows the resulting $\chi^{2}$ calculated from this pruned data set. This $\chi^{2}$ is minimized at $a_{t}=0.689$, very close to the Davies point $a^{*}$.
Figure 7: $\chi^{2}$ as a function of the theoretical value $a_{t}$ calculated from a pruned data set of $151$ observed remnant spin data points. The $a_{t}$ value that minimizes $\chi^{2}$ is $a_{t}=0.689$, indicated with the red arrow, and very near $a^{*}=0.68125$.